OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution. (x - 3) /x = (x - 3) / (x - 6)

OpenStudy (anonymous):

(x-3)/x=(x-3)/(x-6) cross multiply (x-3)(x-6)=x(x-3) now FOIL/distribute (x^2-6x-3x+18)=x^2-3x subtract x^2-3x 18-6x=0 now solve for x -6x=-18 x=3 thats it